OpenStudy (anonymous):

An aluminum plate is made according to the following dimensions. The outer radius is 10 inches and the inner radius is 3 inches. Use p = 3.14 as an approximate value. (a) How many square inches in area is the aluminum plate (the shaded region)? (b) If the cost to construct the plate is $7.50 per square inch, what is the construction cost?

OpenStudy (radar):

\[(\pi)10^{2}-(\pi)3^{2}\]

OpenStudy (radar):

that is the area of the plate in sq in. now multiply cost of $7.50 per square in.

OpenStudy (anonymous):

Multiply it by what?

OpenStudy (radar):

Calculate area as shown in my first response (100pi-9pi) then multiply your answer by $7.50

OpenStudy (anonymous):

I still don't know what my answer is.. I think I'm doing something wrong?

OpenStudy (radar):

Use your calculator and solve the area. Do this Pi(100-9)=91pi. Multiply Pi by 91 the results is the area. Multiplly the area by $7.50 that will be your cost.